# LLM Uncertainty Quantification through Directional Entailment Graph and Claim Level Response Augmentation

### v1

@@ 2 Related Work
The first branch of research is solving the UQ in LLMs by inducing the LLMs to output their uncertainty along with the response Kadavath et al. (2022); Lin et al. (2022); Mielke et al. (2020); Tian et al. (2023). Most of the literature above requires the token-level probabilities of LLMs to train (or fine-tune) and predict the uncertainty. This is a straightforward solution while having full access to model structure and weights, but it can be unwieldy as of time-consuming and resource-tense. Another method Kuhn et al. (2023) estimates LLMs uncertainty directly from response level semantic entropy, yet still requires the token-related probability values as input, which is hard to access given black-box or commercial language models.
In consideration of fast and light-weight evaluation, some researchers propose to solve the UQ by treating the LLMs as black-box and analyzing the consistency in the response semantics structure. Lin et al. (2023) first analyzes the UQ by text responses, treating the sum of eigenvalues from the graph Laplacian as the uncertainty indicator. Chen and Mueller (2023) identify unreliable or speculative answers by computing a confidence score for its generated outputs. However, they solely analyze the UQ from semantics, and  Lin et al. (2023) take the average of entailment probability from two directions to construct a similarity matrix, while in this paper, we find the claims together with the semantics information, better contribute to more comprehensive uncertainty quantification, and the directional logic, is not negligible in nuance analysis of response intrinsic structures.
@@ 3.2 Black Box Evaluation
Given a list of responses $R=\{r_{1},r_{2},\ldots,r_{n}\}$, the pairwise similarity score $s_{ij}$ between responses $r_{i}$ and $r_{j}$ can be calculated using a general similarity function as: $s_{ij}=\text{sim}(r_{i},r_{j})$ the current explorations on matrix $\mathcal{S}$ is mainly based on symmetric similarity property calculations such as Jaccard similarity or worldVector similarity, implying the $s_{ij}=s_{ji}$ in $\mathcal{S}$. Then this condition guarantees the use of Normalized Laplacian to understand the hidden structure in the responses space Lin et al. (2023):
@@ 3.3 Discussion
Since the white box evaluation places a strict requirement on the original model, we analyze from the perspective of the black box evaluation in this paper.
First, the current black box evaluation make assumption that $s_{ij}=s_{ji}$, however, in the actual knowledge representation logic, this neglected the directional information of two responses: If proposition $A$ entails proposition $B$ (denoted as $A\vdash B$), it means that if $A$ is true, then $B$ must also be true. This is a one-way relationship. Importantly, this relationship is not necessarily symmetric; that is, $A\vdash B$ does not imply $B\vdash A$. So the construction of a symmetric matrix broke this rule, which will lead to the loss of directional information from the original response set. In this paper, we propose to reconstruct the response relationship from a directional graph and provide a Random Walk Laplacian uncertainty evaluation method to better fit the asymmetric property of the constructed graph.
@@ 4.2 Enhance the Graph with Text Similarity
Let $\mathcal{S}=[s_{ij}]$ and $\mathcal{T}=[t_{ij}]$ represent the entailment and text similarity matrix, respectively. We define the weight of the edge from node $i$ to node $j$ in the graph $\mathcal{G}_{d}$ as: $w_{ij}=s_{ij}+t_{ij}$. The adjacency matrix $A_{ij}$ of the graph $G$ is can be updated with weights $w_{ij}$.
@@ 5 Claim Based Response Augmentation
To realize it, the key is to first identify the claim atoms in a response $r_{i}$, this step can be achieved with the basic understanding ability of context, we verified that Llama-3 is adequate for this task and used it in the claim extraction. Then to extend extracted claims by recalling the questions, this helps to align the claim descriptions with questions. And at last, combine the augmented claims into a more comprehensive answer $r_{i}^{aug}$ as following:
@@ 6 Experiment
In this section, we design experiment to empirically demonstrate the effectiveness of our proposal in uncertainty evaluation for LLMs. Please note that if without extra declarition, D-UE mean the directional entailment uncertainty on augmented response sets. We intend to investigate the following research questions:
RQ1: Can D-UE improve the uncertainty evaluation layered on existing methods that has no consideration of directional entailment logic?
@@ 6.1 Experiment Setups
As discussed in the paper Lin et al. (2023), there exists the limitation of commonly adopted AUROC that it is very sensitive to imbalanced scenarios (likely to provide over-optimistic evaluation). Area Under Accuracy Rejection Curve (AUARC), is an alternative metric that can better reflect the evaluation performance, the calculation is shown in Appendix B, we use these two as a complementary evaluation indicator.
@@ 6.2 Experiment Result and Analysis
RQ3: We conducted ablation experiments to understand the contribution of the directional entailment uncertainty measure, the claim-based augmentation. Due to the limited page, three representative baselines on two metrics are shown in Fig. 6. We could observe from the baseline of NumSet that, the basic NumSet uncertainty measurement is sensitive to the augmentation by showing improvement in the augmented version (purple bar) over the basic version (blue one), both on AUROC and AUARC. But compared to the improvement brought by claim augmentation, the ‘Ours’ method(D-UE) makes a larger contribution to the general evaluation performance. This attribute to the advantage of directional entailment logic is mined with Random Walk Laplacian on the directional graph. We leave further exploration for the future on how to better combine the semantics uncertainty and directional entailment-based uncertainty and we believe there is still a potential space of improvement based on the current proposed direction.
@@ Appendix B The Evaluation Metric
The AUROC is calculated by plotting the accuracy of accepted predictions against the rejection rate, and then computing the area under this curve.
@@ E.2 The Prompt Used for Response Evaluation
<<INST>><<SYS>> You are given a question, a reference (ground truth) answer, and an actual answer in each round of the task. Task Rate the level of consistency between the actual answer to the reference answer in each question.
@@ E.3 The Prompt Used for Response Claim Augmentation
<<INST>><<SYS>> You are given two pieces of text identified as a question and response claim. A claim is a statement that asserts something as true or false, which can be verified by humans. Task Your task is to first understand every claim stated in the provided text. Then, augment the claims by considering the question being asked, complete the sentence if any part is missing, and resolve any coreference (pronouns or other referring expressions) in the claim for clarity.

### v2

@@ 2 Related Work
The first branch of research is solving the UQ in LLMs by inducing the LLMs to output their uncertainty along with the response Kadavath et al. (2022); Lin et al. (2022); Mielke et al. (2020); Tian et al. (2023). Most of the literature above requires the token-level probabilities of LLMs to train (or fine-tune) and predict the uncertainty. This is a straightforward solution while having full access to model structure and weights, but it can be unwieldy as it is time-consuming and resource-tense. Another method Kuhn et al. (2023) estimates LLMs uncertainty directly from response level semantic entropy, yet still requires the token-related probability values as input, which is hard to access given black-box or commercial language models.
In consideration of fast and lightweight evaluation, some researchers propose to solve the UQ by treating the LLMs as black-box and analyzing the consistency in the response semantics structure. Lin et al. (2023) first analyzes the UQ by text responses, treating the sum of eigenvalues from the graph Laplacian as the uncertainty indicator. Chen and Mueller (2023) identify unreliable or speculative answers by computing a confidence score for its generated outputs. However, they solely analyze the UQ from semantics, and  Lin et al. (2023) take the average of entailment probability from two directions to construct a similarity matrix, while in this paper, we find the claims together with the semantics information, better contribute to more comprehensive uncertainty quantification, and the directional logic, is not negligible in nuance analysis of response intrinsic structures.
@@ 3.2 Black Box Evaluation
Given a list of responses $R=\{r_{1},r_{2},\ldots,r_{n}\}$, the pairwise similarity score $s_{ij}$ between responses $r_{i}$ and $r_{j}$ can be calculated using a general similarity function as: $s_{ij}=\text{sim}(r_{i},r_{j})$ the current explorations on matrix $\mathcal{S}$ are mainly based on symmetric similarity property calculations such as Jaccard similarity or worldVector similarity, implying the $s_{ij}=s_{ji}$ in $\mathcal{S}$. Then this condition guarantees the use of Normalized Laplacian to understand the hidden structure in the responses space Lin et al. (2023):
@@ 3.3 Discussion
Since the white box evaluation places a strict requirement on the original model, we analyze it from the perspective of the black box evaluation in this paper.
First, the current black box evaluation makes assumption that $s_{ij}=s_{ji}$, however, in the actual knowledge representation logic, this neglected the directional information of two responses: If proposition $A$ entails proposition $B$ (denoted as $A\vdash B$), it means that if $A$ is true, then $B$ must also be true. This is a one-way relationship. Importantly, this relationship is not necessarily symmetric; that is, $A\vdash B$ does not imply $B\vdash A$. So the construction of a symmetric matrix broke this rule, which will lead to the loss of directional information from the original response set. In this paper, we propose to reconstruct the response relationship from a directional graph and provide a Random Walk Laplacian uncertainty evaluation method to better fit the asymmetric property of the constructed graph.
@@ 4.2 Enhance the Graph with Text Similarity
Let $\mathcal{S}=[s_{ij}]$ and $\mathcal{T}=[t_{ij}]$ represent the entailment and text similarity matrix, respectively. We define the weight of the edge from node $i$ to node $j$ in the graph $\mathcal{G}_{d}$ as: $w_{ij}=s_{ij}+t_{ij}$. The adjacency matrix $A_{ij}$ of the graph $G$ can be updated with weights $w_{ij}$.
@@ 5 Claim Based Response Augmentation
To realize it, the key is to first identify the claim atoms in a response $r_{i}$, this step can be achieved with the basic understanding ability of context, we verified that Llama-3 is adequate for this task and used it in the claim extraction. Then to extend extracted claims by recalling the questions, this helps to align the claim descriptions with questions. And at last, combine the augmented claims into a more comprehensive answer $r_{i}^{aug}$ as follows:
@@ 6 Experiment
In this section, we design experiments to empirically demonstrate the effectiveness of our proposal in uncertainty evaluation for LLMs. Please note that if without extra declarition, D-UE mean the directional entailment uncertainty on augmented response sets. We intend to investigate the following research questions:
RQ1: Can D-UE improve the uncertainty evaluation layered on existing methods that have no consideration of directional entailment logic?
@@ 6.1 Experiment Setups
As discussed in the paper Lin et al. (2023), there exists the limitation of commonly adopted AUROC that it is very sensitive to imbalanced scenarios (likely to provide over-optimistic evaluation). Area Under Accuracy Rejection Curve (AUARC), is an alternative metric that can better reflect the evaluation performance, the calculation is shown in Appendix B, and we use these two as a complementary evaluation indicator.
@@ 6.2 Experiment Result and Analysis
RQ3: We conducted ablation experiments to understand the contribution of the directional entailment uncertainty measure, the claim-based augmentation. Due to the limited page, three representative baselines on two metrics are shown in Fig. 6. We could observe from the baseline of NumSet that, the basic NumSet uncertainty measurement is sensitive to the augmentation by showing improvement in the augmented version (purple bar) over the basic version (blue one), both on AUROC and AUARC. But compared to the improvement brought by claim augmentation, the ‘Ours’ method(D-UE) makes a larger contribution to the general evaluation performance. This attribute to the advantage of directional entailment logic is mined with Random Walk Laplacian on the directional graph. We leave further exploration for the future on how to better combine the semantics uncertainty and directional entailment-based uncertainty and we believe there is still a potential space for improvement based on the current proposed direction.
@@ Appendix B The Evaluation Metric
Apart from the commonly adopted AUROC that computes the area under the ROC curve, we also employ AUARC for this paper as an evaluation metric for a more comprehensive and sensitive analysis. The AUARC is calculated by plotting the accuracy of accepted predictions against the rejection rate, and then computing the area under this curve.
@@ E.2 The Prompt Used for Response Evaluation
<<INST>><<SYS>> You are given a question, a reference (ground truth) answer, and an actual answer in each round of the task. [Task] Rate the level of consistency between the actual answer to the reference answer in each question.
@@ E.3 The Prompt Used for Response Claim Augmentation
<<INST>><<SYS>> You are given two pieces of text identified as a question and response claim. A claim is a statement that asserts something as true or false, which can be verified by humans. [Task] Your task is to first understand every claim stated in the provided text. Then, augment the claims by considering the question being asked, complete the sentence if any part is missing, and resolve any coreference (pronouns or other referring expressions) in the claim for clarity.